in how many ways can a committee of three men and four woman be formed from a group of twelve men and twelve woman


1
Expert's answer
2021-10-22T14:28:33-0400

Number of ways to of forming 3 men from 12 men "= \frac{12!}{3!(12-3)!} = \frac{10 \times 11 \times 12}{2 \times 3} = 220"

Number of ways to of forming 4 woman from 12 woman "= \frac{12!}{4!(12-4)!} = \frac{9 \times 10 \times 11 \times 12}{2 \times 3 \times 4} = 495"

Number of ways of forming a committee of three men and four woman "= 220 \times 495 = 108900 \;ways"